What is the angle between the arrows of a clock at 3:15? Record your answer
as precisely as possible. (Hint: The arrows are not at exactly the same place.)

Found 2 solutions by stanbon, RAY100:
Answer by stanbon(75887)

What is the angle between the arrows of a clock at 3:15? Record your answer
as precisely as possible. (Hint: The arrows are not at exactly the same place.)
----------------
At 3:15 the small-hand is (1/4) the way from 3 to 4
From 3 to 4 is 360/12 = 30 degrees
(1/4)*30Degrees = 7.5 degrees.
-----------------------------------------
The angle is 7.5 degrees
------------
Cheers,
Stan H.

Answer by RAY100(1637)

assume the 3:15 is 3 hrs and 15 minutes from zero. AM or PM is not given.
The hr hand has rotated once for every hour plus 90 degrees for 15 minutes or
hr = 360 *3 +90 = 1170 degrees.
the minute hand has rotated 60 times for each hour plus 90 degrees for 15 minutes or
minute hand = 3 *60*360 +90= 64,890 degrees
The difference is 63,720